OpenStudy (anonymous):

Find the magnitude & direction of the vector JK given points J(1,0) & K(-2,3).

jimthompson5910 (jim_thompson5910):

First we need to find the vector in <x,y> form Subtract the corresponding coordinates (tail - head) x = -2 - 1 = -3 y = 3 - 0 = 3 the vector JK is <-3, 3> now you will use the formula sqrt(x^2 + y^2) to get the magnitude to get the direction, use arctan(y/x). The angle is in Q2, so you'll have to add 180 degrees

OpenStudy (jdoe0001):

\(\large u=<a,b>\qquad \begin{cases} magnitude\implies \sqrt{a^2+b^2}\\ direction\implies tan^{-1}\left( \frac{b}{a} \right) \end{cases}\)

OpenStudy (anonymous):

wait im kinda confused

jimthompson5910 (jim_thompson5910):

where are you stuck?

OpenStudy (anonymous):

nevermind i got it thanks